OpenStudy (anonymous):

Solve by substitute {3x+2y=18 {3x+8y=36

OpenStudy (agreene):

\[3x+2y=18\\3x=18-2y\\x=\frac{18-2y}{3}\\x=\frac{18}3-\frac{2y}{3}\\x=6-\frac{2y}{3}\]plug this in for the second \[3(6-\frac{2y}3)+8y=36\] solve for y, then plug that in for y in the first eqn to find x.

OpenStudy (anonymous):

So y=3?

OpenStudy (mokeira):

\[3x+2y=18\] \[2y=18-3x\] divide by 2 \[y=9-1.5x\] Now take the second equation and substitute all values of y \[3x+8y=36\] \[3x+8(9-1.5x)=36\] \[3x+72-12x=36\] Collect Like terms \[72-36=12x-3x\] \[36=9x\] Simplify \[x=4 \] Go back to the equation where you found y \[y=9-1.5x\] You have x, so you can use it to find y \[y=9-1.5(4)\] \[y=9-6\] \[y=3 \]

hero (hero):

The easiest way to solve it is by using elimination, however, the best way to substitute is to isolate 3x in both equations: 3x = 18 - 2y 3x = 36 - 8y 3x = 3x 18 - 2y = 36 - 8y 8y - 2y = 36 - 18 6y = 18 y = 3